OpenStudy (anonymous):

How do you integrate the quantity (lnx)^5 divided by 5x?

hartnn (hartnn):

@pirateguyvtcta Hi, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) \(\large \int \dfrac{(\ln x)^5}{5x}dx\) you can put u= ln x then, du=... ? so, (dx/x)=....?